OpenStudy (anonymous):

A circle is represented by the equation below: (x - 9)2 + (y + 8)2 = 16 Which statement is true? A. The circle is centered at (−9, 8) and has a radius of 8. B. The circle is centered at (9, −8) and has a diameter of 8. C. The circle is centered at (9, −8) and has a radius of 8. D. The circle is centered at (−9, 8) and has a diameter of 8.

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

Let us first remind ourselves that : General Equation for a circle whose centre is (h,k) is : \((x -h)^2 + (y -k)^2 = r^2\)

mathslover (mathslover):

Compare this equation with the given one in the question. After comparing, what do you get the values for h, k and r ?

mathslover (mathslover):

For convenience , right the equation given in the question as : \((x-9)^2 + (y-(-8))^2 = 16\)

OpenStudy (anonymous):

B

mathslover (mathslover):

@MLMAQHUZU - yes it is B but please don't just give direct answers. For more info. read OpenStudy's Code of Conduct and Privacy Policy

OpenStudy (anonymous):

ohhhhhhhhh sorry guys, how am I suppose to answer?

OpenStudy (anonymous):

@mathlover- thanks man I'm covered